# Anomalous proximity effect under Andreev and Majorana bound states

### v2

@@ Abstract
Keywords:
Majorana Zero Modes Andreev Bound states Anomalous Proximity Effect Odd-Frequency Pairing
@@ 1 Introduction
The search for Majorana bound states (MBSs) in semiconductor-superconductor systems has spurred enormous efforts since their initial prediction Aguado (2017); Lutchyn et al. (2018); Zhang et al. (2019); Frolov et al. (2020); Flensberg et al. (2021); Prada et al. (2020), in great part because they characterize a topological superconducting state Tanaka et al. (2012); Sato and Fujimoto (2016); Sato and Ando (2017); Tanaka et al. (2024) and also due to their promise for realizing topological qubits Sarma et al. (2015); Beenakker (2020); Aguado and Kouwenhoven (2020); Marra (2022). MBSs emerge as zero-energy edge states after a topological phase transition driven by an applied Zeeman field Tanaka et al. (2024), an idea that triggered an impressive amount of theoretical and experimental studies addressing the zero-energy nature of MBSs Aguado (2017); Lutchyn et al. (2018); Zhang et al. (2019); Frolov et al. (2020); Flensberg et al. (2021); Prada et al. (2020); Tanaka et al. (2024). Despite the apparent simple recipe, the detection of MBSs has turned out to be difficult because semiconductor-superconductor systems also host trivial zero-energy Andreev bound states (ABSs) Bagrets and Altland (2012); Pikulin et al. (2012); Kells et al. (2012); Cayao et al. (2015); Prada et al. (2012); Reeg et al. (2018); Awoga et al. (2019); Das Sarma and Pan (2021); Awoga et al. (2022); Cayao and Black-Schaffer (2021); Cayao and Burset (2021); Marra and Nigro (2022); de Mendonça et al. (2023); Baldo et al. (2023); Awoga et al. (2023); Awoga and Cayao (2024); Cayao (2024); Hess et al. (2024); Legg (2025a); Legg et al. (2024); Legg (2025b) that can behave as MBSs. As a result, identifying robust zero-energy states in semiconductor-superconductor systems is only a necessary condition for MBSs Tanaka and Kashiwaya (1995); Kashiwaya et al. (1996); Kashiwaya and Tanaka (2000); Tanaka and Kashiwaya (2004); Law et al. (2009); Flensberg (2010) and, to approach a sufficient condition, the detection protocol must also involve simultaneous measurements of other Majorana properties Prada et al. (2017); Cayao et al. (2017); Clarke (2017); Cayao et al. (2018); Cayao and Black-Schaffer (2021); Pekerten et al. (2022); Vimal and Cayao (2024); Mizushima et al. (2025).
@@
In the mid-2000s, the simultaneous analysis of multiple Majorana properties was shown to be unambiguously revealed by the so-called anomalous proximity effect in a disordered normal metal (DN) attached to a spin-triplet superconductor Tanaka and Kashiwaya (2004); Tanaka et al. (2005a, b); Asano et al. (2006a). In the anomalous proximity effect, zero-energy MBSs emerging at the interface penetrate into the DN region, giving rise to a zero-energy peak (ZEP) in the local density of states (LDOS) Tanaka and Kashiwaya (2004); Tanaka et al. (2005a, b); Asano et al. (2007); Higashitani et al. (2009). In contrast, for a spin-singlet $s$-wave superconductor, the conventional proximity effect is characterized by a zero-energy dip in the LDOS Golubov and Kupriyanov (1988); Belzig et al. (1996). Further studies revealed that the main actor in the anomalous proximity effect is the odd-frequency spin-triplet s-wave pairing induced into the disordered normal region Tanaka and Golubov (2007); Tanaka et al. (2007a, b); Higashitani et al. (2009); Tanaka et al. (2012); Asano and Tanaka (2013); Tanaka et al. (2022); Kokkeler et al. (2023). The s-wave nature of this type of pairing was shown to be crucial for its robustness against scalar disorder. Moreover, the anomalous proximity effect has been also shown to originate other exotic phenomena, such as enhanced Josephson currents at low temperatures Asano et al. (2006a, b), paramagnetic Meissner response Tanaka et al. (2005b); Yokoyama et al. (2011), anomalous surface impedance Asano et al. (2011), and a zero-bias voltage quantized conductance peak Tanaka and Kashiwaya (2004); Tanaka et al. (2005a); Ikegaya et al. (2016); Takagi et al. (2020). This fueled the interest in the anomalous proximity effect, leading to a series of theoretical studies Tanaka et al. (2012); Tamura and Tanaka (2019); Suzuki et al. (2019); Mizushima et al. (2023); Ikegaya et al. (2015, 2016); Ikegaya and Asano (2017); Ikegaya et al. (2018, 2020); Nagae et al. (2025) and culminating in the recent experimental detection of the anomalous proximity effect in CoSi2/TiSi2 heterostructures Chiu et al. (2021, 2023, 2024). In all cases, the anomalous proximity effect appears as a strong signature of unconventional spin-triplet superconductivity, suggesting that it can be also used to identify and understand other unconventional superconducting states. Given that robust zero-energy ABSs in semiconductor-superconductor systems can behave as MBSs, and the emergent superconductivity is unconventional, it is natural to ask if such states produce an anomalous proximity effect which can improve our understanding of Majorana and Andreev physics.
@@ 2 Model and Methods
We consider a one-dimensional semiconductor nanowire with strong spin-orbit coupling and proximate the right part of it to a conventional $s$-wave superconductor. We also assume that the right portion of the unproximated region hosts scalar impurities, effectively realizing a clean normal metal/disordered normal metal/superconductor (CN/DN/S) junction. The junction is subjected to an external Zeeman field along the $x$ direction. The value of the Zeeman field determines whether we are in the topological or trivial regime. We model the system using a one-dimensional tight-binding Hamiltonian on a lattice with lattice constant $a$. The Bogoliubov-de Gennes (BdG) Hamiltonian for the entire system can be written as:
[…]
@@
with $\mu_{N}<\mu_{S}$.
Disorder is introduced in the DN region via a random onsite potential $v_{i}$ drawn from a uniform distribution $[-W,W]$, where $W$ is the scalar onsite disorder strength. Outside the DN region, the system is considered to be clean ($v_{i}=0$ for $i\leq(L_{\rm N}-L_{\rm DN})/a$ and for $i>L_{\rm N}/a$). We refer to $W\lesssim\mu_{S}$ as weak disorder, $W$ comparable to a few times $\mu_{S}$ (e.g., $W\approx(1\sim 3)\mu_{S}$) as moderate disorder, and $W\gg\mu_{S}$ as strong disorder in our discussions.
@@
The topological phase transition in the S region occurs when the condition $B=B_{c}=\sqrt{\Delta^{2}+\mu_{S}^{2}}$ is satisfied, provided that the wire is sufficiently long Tanaka et al. (2024). In the topological phase ($B>B_{c}$), MBSs are expected to form at the ends of the S region.
The system has another phase, the helical phase in the normal region. The helical phase starts at the helical transition point $B=\mu_{N}$ and extends to the region $\mu_{N}<B<B_{c}$. In this phase, the system is in a topologically trivial state and the zero-energy states are expected to be trivial ABSs Cayao et al. (2015); Cayao and Black-Schaffer (2021). The helical phase is characterized by the presence of a zero-energy state bound to the DN/S interface, which is a consequence of the helical nature of the system Ahmed et al. (2024).
@@
where $\hat{}$ denotes a matrix in spin space, $d_{s}$ is the spin-singlet component and $\vec{d}=(d_{x},d_{y},d_{z})^{T}$ are the spin-triplet components of the pair correlations, and we have dropped spatial indices for clarity. It has been shown that only local pairing correlations can survive in the presence of disorder and hence, we only consider local pairing correlations in this paper Tanaka and Golubov (2007). The local pair correlations can be calculated as:
@@
Due to fermionic statistics, pair amplitudes must be overall antisymmetric under particle exchange. For local ($s$-wave) correlations, spatial symmetry is even. Therefore, the spin-singlet component $d_{s}$ must be even in frequency, while the spin-triplet components $d_{x}$, $d_{y}$, and $d_{z}$ must be odd in frequency Bergeret et al. (2005); Tanaka et al. (2012, 2024); Linder and Balatsky (2019); Cayao et al. (2020); Triola et al. (2020). Having outlined the model and methods, we now list the parameters used in this work. We use realistic parameters for our model based on InSb and InAs nanowires Lutchyn et al. (2018). For numerical convenience, we choose a large lattice constant $a=50$ nm, which is five times larger than the value used in the literature. This allows us to use a smaller number of lattice sites in our calculations. The parameters used are: the hopping amplitude $t=1$ meV, the chemical potential in the superconductor $\mu_{S}=0.5$ meV, the chemical potential in the normal region $\mu_{N}=0.1$ meV, the superconducting gap $\Delta=0.25$ meV, and the spin-orbit coupling strength $\alpha=20$ meV nm. The disorder strength $W$ is varied between 00 and $5\mu_{S}$. The length of the S region $L_{\rm S}$ is varied between $L_{\rm S}=20a$ and $L_{\rm S}=100a$), while the total length of the normal region is fixed at $L_{\rm N}=40a$. The length of the disordered region $L_{\rm DN}$ is varied ($L_{DN}=2a,20a,40a$), which determines the length of the clean normal region $L_{\rm CN}=L_{\rm N}-L_{\rm DN}$. The critical Zeeman field for the topological transition is $B_{c}=\sqrt{\Delta^{2}+\mu_{S}^{2}}\approx 0.56$ meV. The broadening parameter is set to $\eta=10^{-3}$. Results involving disorder are averaged over 200 independent of disorder realizations, unless otherwise specified.
@@ 3.1 Low-Energy Spectrum for Short S Regions
In the clean limit $W=0$ (light gray curves in Fig. 2 (a,b,d,e,g,h)), the spectrum exhibits the expected behavior: the bulk superconducting gap closes and reopens around the critical field $B_{c}=\sqrt{\Delta^{2}+\mu_{S}^{2}}\approx 0.56$ meV, marking the topological phase transition. For $B>B_{c}$, zero-energy MBSs appear at the ends of the S region. Due to the finite length $L_{S}=20a$, these MBSs hybridize, leading to energy eigenvalues that oscillate around $E=0$ as $B$ increases. The spectrum in the clean limit has another gap closing at $B=\mu_{N}$ purely due to confinement and helicity in the N regionAhmed et al. (2024), signaling that the system has entered into a trivial helical regime. In the helical regime ($\mu_{N}<B<B_{c}$), ABSs emerge, and their energies oscillate around zero in a similar fashion to the Majorana oscillations in the topological regime.
@@
Figure 4: LDOS for short S region ($L_{\rm S}=20a$) junctions evaluated at the leftmost site of the N region ($i=1$) as a function of energy $E$ and scalar disorder strength $W$ with $L_{\rm DN}=2a$ (a,f), $L_{\rm DN}=20a$ (b,g), and $L_{\rm DN}=40a$ (c,h). Panels (d,i) show line cuts of the LDOS in (a,f,b,g) for $W/\mu_{S}=0,1,2$. Panels (e,j) show the spatial profile of the LDOS at $E\approx 0$ for $W/\mu_{S}=0,1,2$ with $L_{\rm DN}=20a$. Top row (a,b,c,d,e): trivial phase ($B=0.825B_{c}$). Bottom row (f,g,h,i,j): topological phase ($B=1.2B_{c}$). Other parameters used are: $L_{\rm S}=20a$, $L_{\rm N}=40a$, $a=50$ nm, $\mu_{S}=0.5$ meV, $\mu_{N}=0.1$ meV, $\Delta=0.25$ meV, $\alpha=20$ meV nm, $t=1$ meV, and $\eta=0.001$ meV.
@@
Figure 5: Absolute value of the spin-triplet pair correlations $|\vec{d}|$ for short S region ($L_{\rm S}=20a$) junctions evaluated at the leftmost site of the N region ($i=1$) as a function of energy $E$ and scalar disorder strength $W$ with $L_{\rm DN}=2a$ (a,f), $L_{\rm DN}=20a$ (b,g), and $L_{\rm DN}=40a$ (c,h). Panels (d,i) show line cuts of the spin-triplet pair correlations in (a,f,b,g) for $W/\mu_{S}=0,1,2$. Panels (e,j) show the spatial profile of the spin-triplet pair correlations at $E\approx 0$ for $W/\mu_{S}=0,1,2$ with $L_{\rm DN}=20a$. Top row (a,b,c,d,e): trivial phase ($B=0.825B_{c}$). Bottom row (f,g,h,i,j): topological phase ($B=1.2B_{c}$). Other parameters used are: $L_{\rm S}=20a$, $L_{\rm N}=40a$, $a=50$ nm, $\mu_{S}=0.5$ meV, $\mu_{N}=0.1$ meV, $\Delta=0.25$ meV, $\alpha=20$ meV nm, $t=1$ meV, and $\eta=0.001$ meV.
@@
Figure 6: Absolute value of the spin-singlet pair correlation for short S region ($L_{\rm S}=20a$) junctions evaluated at the leftmost site of the N region ($i=1$) as a function of energy $E$ and scalar disorder strength $W$ with $L_{\rm DN}=2a$ (a,f), $L_{\rm DN}=20a$ (b,g), and $L_{\rm DN}=40a$ (c,h). Panels (d,i) show line cuts of the spin-singlet pair correlations in (a,f,b,g) for $W/\mu_{S}=0,1,2$. Panels (e,j) show the spatial profile of the spin-singlet pair correlations at $E\approx 0$ for $W/\mu_{S}=0,1,2$ with $L_{\rm DN}=20a$. Top row (a,b,c,d,e): trivial phase ($B=0.825B_{c}$). Bottom row (f,g,h,i,j): topological phase ($B=1.2B_{c}$). Other parameters used are: $L_{\rm S}=20a$, $L_{\rm N}=40a$, $a=50$ nm, $\mu_{S}=0.5$ meV, $\mu_{N}=0.1$ meV, $\Delta=0.25$ meV, $\alpha=20$ meV nm, $t=1$ meV, and $\eta=0.001$ meV.
@@
The spatial profile of the zero-energy bound states is shown in Fig. 2(c,f,i) for a single disorder realization (solid lines) and for the ensemble average (dashed lines) for $W=\mu_{S}$ and $W=2\mu_{S}$. The top panels show the wavefunction probability density for the trivial zero-energy ABSs at $B=0.825B_{c}$, while the bottom panels show the wavefunction probability density for the MBSs at $B=1.2B_{c}$. We see that for short DN regions (Fig. 2(c)), disorder has little effect on the wavefunction profile of the zero-energy states. In the helical phase, the wavefunction is localized at the DN/S interface with an oscillatory profile in the N region. In the topological phase, the wavefunction is localized at the ends of the S region, with the MBSs having a similar oscillatory profile across the entire N region, in agreement with the clean limit Ahmed et al. (2024). For longer DN regions (Fig. 2(f,i)), the wavefunction profile of the zero-energy states is significantly affected by disorder. While the wavefunctions of both the trivial and topological zero-energy states remain localized at the NS interface and the ends of the S region, respectively, the oscillatory profile in the entire N region is lost. Instead, the wavefunction profile is localized at the DN/S interface with an exponential decay into the DN region. This decay continues even in the clean region. The wavefunction profile of the zero-energy states is also affected by the disorder strength $W$. As disorder increases, the wavefunction profile becomes more localized at the DN/S interface and decays more rapidly into the N region. This is particularly evident in the topological phase, where the wavefunction of the MBSs becomes more localized at the DN/S interface rather than leaking into the N region, as observed in the clean limit.

### v3

@@ 1 Introduction
The search for Majorana bound states (MBSs) in semiconductor-superconductor systems has spurred enormous efforts since their initial prediction Aguadoreview17 (1, 2, 3, 4, 5, 6), in great part because they characterize a topological superconducting state tanaka2012symmetry (7, 8, 9, 10) and also due to their promise for realizing topological qubits sarma2015majorana (11, 12, 13, 14). MBSs emerge as zero-energy edge states after a topological phase transition driven by an applied Zeeman field tanaka2024theory (10), an idea that triggered an impressive amount of theoretical and experimental studies addressing the zero-energy nature of MBSs Aguadoreview17 (1, 2, 3, 4, 5, 6, 10). Despite the apparent simple recipe, the detection of MBSs has turned out to be difficult because semiconductor-superconductor systems also host trivial zero-energy Andreev bound states (ABSs) Bagrets:PRL12 (15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35) that can behave as MBSs. As a result, identifying robust zero-energy states in semiconductor-superconductor systems is only a necessary condition for MBSs TK95 (36, 37, 38, 39, 40, 41) and, to approach a sufficient condition, the detection protocol must also involve simultaneous measurements of other Majorana properties PhysRevB.96.085418 (42, 43, 44, 45, 24, 46, 47, 48, 49).
@@
In the mid-2000s, the simultaneous analysis of multiple Majorana properties was shown to be unambiguously revealed by the so-called anomalous proximity effect in a disordered normal metal (DN) attached to a spin-triplet superconductor Proximityp (39, 50, 51, 52). In the anomalous proximity effect, zero-energy MBSs emerging at the interface penetrate into the DN region, giving rise to a zero-energy peak (ZEP) in the local density of states (LDOS) Proximityp (39, 50, 51, 53, 54). In contrast, for a spin-singlet $s$-wave superconductor, the conventional proximity effect is characterized by a zero-energy dip in the LDOS golubov1988theoretical (55, 56). Further studies revealed that the main actor in the anomalous proximity effect is the odd-frequency spin-triplet s-wave pairing induced into the disordered normal region odd1 (57, 58, 59, 54, 7, 60, 61, 62). The s-wave nature of this type of pairing was shown to be crucial for its robustness against scalar disorder. Moreover, the anomalous proximity effect has been also shown to originate other exotic phenomena, such as enhanced Josephson currents at low temperatures PhysRevLett.96.097007 (52, 63), paramagnetic Meissner response PhysRevB.72.140503 (51, 64), anomalous surface impedance PhysRevLett.107.087001 (65), and a zero-bias voltage quantized conductance peak Proximityp (39, 50, 66, 67). This fueled the interest in the anomalous proximity effect, leading to a series of theoretical studies tanaka2012symmetry (7, 68, 69, 70, 71, 66, 72, 73, 74, 75) and culminating in the recent experimental detection of the anomalous proximity effect in CoSi2/TiSi2 heterostructures doi:10.1126/sciadv.abg6569 (76, 77, 78). In all cases, the anomalous proximity effect appears as a strong signature of unconventional spin-triplet superconductivity, suggesting that it can be also used to identify and understand other unconventional superconducting states.  [^1] Other systems expected to give an anomalous proximity effect are those hosting non-Hermitian effects, such as those reported in Refs. xbj1-hfyf (79, 31, 80, 81, 82, 83, 84) Given that robust zero-energy ABSs in semiconductor-superconductor systems can behave as MBSs, and the emergent superconductivity is unconventional, it is natural to ask if such states produce an anomalous proximity effect which can improve our understanding of Majorana and Andreev physics.
@@ 2 Model and Methods
We consider a one-dimensional semiconductor nanowire with strong spin-orbit coupling. The right segment of the wire is placed on top of a conventional s-wave superconductor. Due to the proximity effect, a superconducting pairing gap $\Delta$ is induced. We also assume that the right portion of the unproximated region hosts scalar impurities, effectively realizing a clean normal metal/disordered normal metal/superconductor (CN/DN/S) junction. The junction is subjected to an external Zeeman field along the $x$ direction. The value of the Zeeman field determines whether we are in the topological or trivial regime. We model the system using a one-dimensional tight-binding Hamiltonian on a lattice with lattice constant $a$. The Bogoliubov-de Gennes (BdG) Hamiltonian for the entire system can be written as:
$H=\sum_{i}\Psi_{i}^{\dagger}\mathcal{H}_{ii}\Psi_{i}+\sum_{\langle i,j\rangle}\left(\Psi_{i}^{\dagger}\mathcal{H}_{ij}\Psi_{j}+\text{H.c.}\right),$ (1)
@@
with $\mu_{N}<\mu_{S}$. We note that our model describes a system in thermodynamic equilibrium. Hence, no net particle current flows despite the chemical potential difference.
Disorder is introduced in the DN region via a random onsite potential $v_{i}$ drawn from a uniform distribution [W,W][-W,W], where $W$ is the scalar onsite disorder strength. Experimentally, such disorder could be realized via ion irradiation or by controlling defects during growth nam2005disorder (85, 86, 87, 88). Outside the DN region, the system is considered to be clean ($v_{i}=0$ for $i\leq(L_{\rm N}-L_{\rm DN})/a$ and for $i>L_{\rm N}/a$). We refer to $W\lesssim\mu_{S}$ as weak disorder, $W$ comparable to a few times $\mu_{S}$ (e.g., $W\approx(1\sim 3)\mu_{S}$) as moderate disorder, and $W\gg\mu_{S}$ as strong disorder in our discussions. We use $\mu_{S}$ as the reference energy scale as it sets the Fermi energy in the superconducting region. Since the topological superconductivity and the associated MBSs originate from the S region, $\mu_{S}$ is the natural energy scale for the quasiparticles involved in the proximity effect. Comparing the disorder potential strength $W$ to $\mu_{S}$ gives a measure of how disruptive the scattering is to these quasiparticles.
@@
The topological phase transition in the S region occurs when the condition $B=B_{c}=\sqrt{\Delta^{2}+\mu_{S}^{2}}$ is satisfied, provided that the wire is sufficiently long tanaka2024theory (10), meaning its length $L_{S}$ is much larger than twice the Majorana localization length ($L_{S}\gg 2\xi_{\rm MBS}$) to suppress the hybridization of end states. In the topological phase ($B>B_{c}$), MBSs are expected to form at the ends of the S region. [^2] Majorana states can be characterized by nontrivial topological invariants v6ht-jq3l (89, 30, 90, 91, 92, 93, 94) as well as by unique features of their odd-frequency superconducting pairing tanaka2012symmetry (7, 10, 95).
The system has another phase, the helical phase in the normal region. The helical phase starts at the helical transition point $B=\mu_{N}$ and extends to the region $\mu_{N}<B<B_{c}$. In this phase, the system is in a topologically trivial state and the zero-energy states are expected to be trivial ABSs PhysRevB.91.024514 (18, 24). The helical phase is characterized by the presence of a zero-energy state bound to the DN/S interface, which is a consequence of the helical nature of the system ahmed2025odd (96).
@@
where $\hat{}$ denotes a matrix in spin space, $d_{s}$ is the spin-singlet component and $\vec{d}=(d_{x},d_{y},d_{z})^{T}$ are the spin-triplet components of the pair correlations, and we have dropped spatial indices for clarity. While the topological superconductivity in the clean S region is known to be of an effective p-wave nature, pair correlations with non-zero angular momentum are fragile against the scalar disorder present in the DN region. It has been shown that only local s-wave pairing correlations can survive in the presence of disorder and hence, we only consider local pairing correlations in this paper odd1 (57). The local pair correlations can be calculated as:
@@
Due to fermionic statistics, pair amplitudes must be overall antisymmetric under particle exchange. For local ($s$-wave) correlations, spatial symmetry is even. Therefore, the spin-singlet component $d_{s}$ must be even in frequency, while the spin-triplet components $d_{x}$, $d_{y}$, and $d_{z}$ must be odd in frequency RevModPhys.77.1321 (97, 7, 10, 98, 95, 99). Having outlined the model and methods, we now list the parameters used in this work. We use realistic parameters for our model based on InSb and InAs nanowires lutchyn2018majorana (2). For numerical convenience, we choose a large lattice constant $a=50$ nm, which is five times larger than the value used in the literature. This allows us to use a smaller number of lattice sites in our calculations. The parameters used are: the hopping amplitude $t=1$ meV, the chemical potential in the superconductor $\mu_{S}=0.5$ meV, the chemical potential in the normal region $\mu_{N}=0.1$ meV, the superconducting gap $\Delta=0.25$ meV, and the spin-orbit coupling strength $\alpha=20$ meV nm. The disorder strength $W$ is varied between 0 and $5\mu_{S}$. The length of the S region $L_{\rm S}$ is varied between $L_{\rm S}=20a$ and $L_{\rm S}=100a$), while the total length of the normal region is fixed at $L_{\rm N}=40a$. The length of the disordered region $L_{\rm DN}$ is varied ($L_{DN}=2a,20a,40a$), which determines the length of the clean normal region $L_{\rm CN}=L_{\rm N}-L_{\rm DN}$. The total length of the normal region is fixed to ensure constant confinement conditions, which is necessary for the formation of confinement-induced helical Andreev bound states in the trivial regime ahmed2025odd (96). By varying $L_{\rm DN}$ within this fixed length, we can isolate the effect of disorder from the effects of changing confinement. The critical Zeeman field for the topological transition is $B_{c}=\sqrt{\Delta^{2}+\mu_{S}^{2}}\approx 0.56$ meV. The broadening parameter is set to $\eta=10^{-3}$. Results involving disorder are averaged over 200 independent of disorder realizations, unless otherwise specified.
@@ 3.1 Low-Energy Spectrum for Short S Regions
In the clean limit $W=0$ (light gray curves in Fig. 2 (a,b,d,e,g,h)), the spectrum exhibits the expected behavior: the bulk superconducting gap closes and reopens around the critical field $B_{c}=\sqrt{\Delta^{2}+\mu_{S}^{2}}\approx 0.56$ meV, marking the topological phase transition. For $B>B_{c}$, zero-energy MBSs appear at the ends of the S region. Due to the finite length $L_{S}=20a$, these MBSs hybridize, leading to energy eigenvalues that oscillate around $E=0$ as $B$ increases. The spectrum in the clean limit has another gap closing at $B=\mu_{N}$ purely due to confinement and helicity in the N regionahmed2025odd (96), signaling that the system has entered into a trivial helical regime. In the helical regime ($\mu_{N}<B<B_{c}$), ABSs emerge, and their energies oscillate around zero in a similar fashion to the Majorana oscillations in the topological regime.
@@
Figure 4: Ensemble-averaged LDOS for short S region ($L_{\rm S}=20a$) junctions evaluated at the leftmost site of the N region ($i=1$) as a function of energy $E$ and scalar disorder strength $W$ with $L_{\rm DN}=2a$ (a,f), $L_{\rm DN}=20a$ (b,g), and $L_{\rm DN}=40a$ (c,h). Panels (d,i) show line cuts of the LDOS in (a,f,b,g) for $W/\mu_{S}=0,1,2$. Panels (e,j) show the spatial profile of the LDOS at $E\approx 0$ for $W/\mu_{S}=0,1,2$ with $L_{\rm DN}=20a$. Top row (a,b,c,d,e): trivial phase ($B=0.825B_{c}$). Bottom row (f,g,h,i,j): topological phase ($B=1.2B_{c}$). Other parameters used are: $L_{\rm S}=20a$, $L_{\rm N}=40a$, $a=50$ nm, $\mu_{S}=0.5$ meV, $\mu_{N}=0.1$ meV, $\Delta=0.25$ meV, $\alpha=20$ meV nm, $t=1$ meV, and $\eta=0.001$ meV.
@@
Figure 5: Ensemble-averaged Absolute value of the spin-triplet pair correlations $|\vec{d}|$ for short S region ($L_{\rm S}=20a$) junctions evaluated at the leftmost site of the N region ($i=1$) as a function of energy $E$ and scalar disorder strength $W$ with $L_{\rm DN}=2a$ (a,f), $L_{\rm DN}=20a$ (b,g), and $L_{\rm DN}=40a$ (c,h). Panels (d,i) show line cuts of the spin-triplet pair correlations in (a,f,b,g) for $W/\mu_{S}=0,1,2$. Panels (e,j) show the spatial profile of the spin-triplet pair correlations at $E\approx 0$ for $W/\mu_{S}=0,1,2$ with $L_{\rm DN}=20a$. Top row (a,b,c,d,e): trivial phase ($B=0.825B_{c}$). Bottom row (f,g,h,i,j): topological phase ($B=1.2B_{c}$). Other parameters used are: $L_{\rm S}=20a$, $L_{\rm N}=40a$, $a=50$ nm, $\mu_{S}=0.5$ meV, $\mu_{N}=0.1$ meV, $\Delta=0.25$ meV, $\alpha=20$ meV nm, $t=1$ meV, and $\eta=0.001$ meV.
@@
Figure 6: Ensemble-averaged Absolute value of the spin-singlet pair correlation for short S region ($L_{\rm S}=20a$) junctions evaluated at the leftmost site of the N region ($i=1$) as a function of energy $E$ and scalar disorder strength $W$ with $L_{\rm DN}=2a$ (a,f), $L_{\rm DN}=20a$ (b,g), and $L_{\rm DN}=40a$ (c,h). Panels (d,i) show line cuts of the spin-singlet pair correlations in (a,f,b,g) for $W/\mu_{S}=0,1,2$. Panels (e,j) show the spatial profile of the spin-singlet pair correlations at $E\approx 0$ for $W/\mu_{S}=0,1,2$ with $L_{\rm DN}=20a$. Top row (a,b,c,d,e): trivial phase ($B=0.825B_{c}$). Bottom row (f,g,h,i,j): topological phase ($B=1.2B_{c}$). Other parameters used are: $L_{\rm S}=20a$, $L_{\rm N}=40a$, $a=50$ nm, $\mu_{S}=0.5$ meV, $\mu_{N}=0.1$ meV, $\Delta=0.25$ meV, $\alpha=20$ meV nm, $t=1$ meV, and $\eta=0.001$ meV.
@@
The spatial profile of the zero-energy bound states is shown in Fig. 2(c,f,i) for a single disorder realization (solid lines) and for the ensemble average (dashed lines) for $W=\mu_{S}$ and $W=2\mu_{S}$. The top panels show the wavefunction probability density for the trivial zero-energy ABSs at $B=0.825B_{c}$, while the bottom panels show the wavefunction probability density for the MBSs at $B=1.2B_{c}$. We see that for short DN regions (Fig. 2(c)), disorder has little effect on the wavefunction profile of the zero-energy states. In the helical phase, the wavefunction is localized at the DN/S interface with an oscillatory profile in the N region. In the topological phase, the wavefunction is localized at the ends of the S region, with the MBSs having a similar oscillatory profile across the entire N region, in agreement with the clean limit ahmed2025odd (96). For longer DN regions (Fig. 2(f,i)), the wavefunction profile of the zero-energy states is significantly affected by disorder. While the wavefunctions of both the trivial and topological zero-energy states remain localized at the NS interface and the ends of the S region, respectively, the oscillatory profile in the entire N region is lost. Instead, the wavefunction profile is localized at the DN/S interface with an exponential decay into the DN region. This decay continues even in the clean region. The wavefunction profile of the zero-energy states is also affected by the disorder strength $W$. As disorder increases, the wavefunction profile becomes more localized at the DN/S interface and decays more rapidly into the N region. This is particularly evident in the topological phase, where the wavefunction of the MBSs becomes more localized at the DN/S interface rather than leaking into the N region, as observed in the clean limit.